Find the orthogonal trajectories of the family of parabolas x=cy²
1
Expert's answer
2019-08-30T09:54:22-0400

Let "(x, \quad y=Y)" be the intersection point of the curves

"y=\sqrt{\frac xc} \quad \perp \quad Y=Y(x), \quad x>0, \quad y>0."

Then

"Y'=-\frac{1}{y'}=-2\sqrt{cx}."

But "c=\frac{x}{y^2}" . Therefore

"Y'=-\frac{2x}{Y} \quad \text{or} \quad Y^2+2x^2=C."

We set the family of ellipses

"\frac{x^2}{a^2}+\frac{y^2}{2a^2}=1."